OpenStudy (anonymous):

3^x-2 = 4^2x+4 can you solve this in ln form?

OpenStudy (anonymous):

take logs of both sides

OpenStudy (anonymous):

wait then when do you use natural logs?

OpenStudy (anonymous):

@Moyo30

OpenStudy (anonymous):

In is natural log so take In of both sides

OpenStudy (anonymous):

okay thats my first step...then what is next?

OpenStudy (anonymous):

what do u get?

OpenStudy (anonymous):

if you take the log of both sides?

OpenStudy (anonymous):

i got \[(x-2)\ln3=(2x+4)\ln4\]

OpenStudy (anonymous):

@Moyo30

OpenStudy (anonymous):

collect like terms

OpenStudy (anonymous):

but dont i have to convert to logs?

OpenStudy (anonymous):

u have used the logs to get to that stage but u need to simplify further

OpenStudy (anonymous):

okay so i got x (ln3-2ln4) = 4ln4+2ln3

OpenStudy (anonymous):

then further simplify to solve for x

OpenStudy (anonymous):

x=(4ln(4) + 2ln3)) / (ln(3) - 2ln(4))

OpenStudy (tkhunny):

You must learn to write more clearly. What you have written 3^x-2 = 4^2x+4 means this \(3^{x} - 2 = \left(4^{2}\cdot x\right) + 4\) If you mean something else, please use sufficient parentheses to clarify. Remember that the Order of Operations ALWAYS works!

OpenStudy (anonymous):

so i got this http://bamboodock.wacom.com/doodler/9459161b-e500-4329-9818-c0bf0d611c8a

OpenStudy (anonymous):

so do i turn into log after this?

OpenStudy (tkhunny):

You can continue to simplfy. Use your log rules.

OpenStudy (anonymous):

i am so lost ><

OpenStudy (tkhunny):

It does not matter if you use log or ln. In this case, it doesn't matter what base you are using. So far, you have done nothing that requires any particular base.

OpenStudy (anonymous):

okay well why is ln and log answers are different?

OpenStudy (tkhunny):

Three possibilities come to mind: 1) They aren't. 2) It is made clear which base to use and you are looking at the wrong one. 3) You are missing something. How shall we determine which it is?

OpenStudy (anonymous):

ln is to log to base e (e is a special number) approximate value of e = 2.718 log is base 10

OpenStudy (anonymous):

okay but the book gave an example of solving log equations and it was using through ln and this was the example http://bamboodock.wacom.com/doodler/e2ff45a5-3ae8-477a-9464-d3bee76dd49b so i thought you solve the equation through ln but there was another problem but the answer book said to use by log and i was lost from there

OpenStudy (tkhunny):

\(3^{x-2} = 4^{2x+4}\) Introduce logs base 10 \((x-2)\cdot\log_{10}(3) = (2x+4)\cdot\log_{10}(4)\) Introduce logs base e Choose from \((x-2)\cdot\log_{e}(3) = (2x+4)\cdot\log_{e}(4)\) \((x-2)\cdot\ln(3) = (2x+4)\cdot\ln(4)\) \((x-2)\cdot\log(3) = (2x+4)\cdot\log(4)\) Introduce logs base 2 \((x-2)\cdot\log_{2}(3) = (2x+4)\cdot\log_{2}(4)\) It does not matter, so long as the context is clear and you remain consistent.

OpenStudy (anonymous):

OMG THANK YOU SO MUCH! this made it so clear!

OpenStudy (anonymous):

thank you thank you thank you!!!